OpenStudy (anonymous):

how would you solve this? (6e^x^2)-e^(7-x^2)=0

OpenStudy (anonymous):

move 2nd term to the right \[6 e^{x^2}= e^{7-x^2} \] then take natural log on both sides: \[\ln(6)+x^2=7-x^2\] so \[x^2=\frac{7-\ln(6)}{2}\] and take sqrt root.... done!

OpenStudy (anonymous):

OMG thanks. I didn't think to take the natural log of 6. Thanks =)